OpenStudy (anonymous):

how would you find the solutions to x^3-x^2-4x+4=0?

OpenStudy (beginnersmind):

You could guess a root. In this case x=1 works, so you can factor x-1 out. Then divide and solve the quadratic the usual way.

OpenStudy (anonymous):

i know the answers... there -2,1 and 2. i know the first step is to factor x^3-x^2-4x+4 when you factor that you get (x-2)(x-1)(x-2)

OpenStudy (anonymous):

i wanna know how to get x-2)(x-1)(x-2)

OpenStudy (anonymous):

how do you factor it

OpenStudy (beginnersmind):

Well, there's a formula, like for quadratics. But believe it or not, it's usual to just guess at some solution x and use long division.

OpenStudy (anonymous):

how will that help me factor?

OpenStudy (anonymous):

Do grouping first. (x^3-x^2) (4x+4) then look what their common factor is.

OpenStudy (anonymous):

i know "x" is

OpenStudy (beginnersmind):

If you have a solution x1 then x-x1 divides the original polynomial. You divide the original polynomial by x-x1 and go on from there. Can you divide polynomials?

OpenStudy (anonymous):

yes i can. http://www.wolframalpha.com/input/?i=x^3-x^2-4x%2B4%3D0#socialwelcome these are the steps, i just wanna know how to factor...

OpenStudy (beginnersmind):

You mean how to factor after you know x=1 is a root or how to find x=1 in the first place?

OpenStudy (anonymous):

u might have to scroll down next to answer and click on step by step solution

OpenStudy (anonymous):

i just wanna know how to factor x^3-x^2-4x+4

OpenStudy (anonymous):

(x^3 - x^2) (4x +4) -x^2 (x+1) 4 (x+1) (-x^2+4)(x+1)

OpenStudy (anonymous):

is this it? If you are factoring wit those numbers, i think you use grouping, right?

OpenStudy (anonymous):

thanks but ur factoring (x^3 - x^2) (4x +4) i need x^3-x^2-4x+4

OpenStudy (anonymous):

you group them

OpenStudy (anonymous):

then you find gcf

OpenStudy (anonymous):

OH! Okay I think i got it now (x^3-x^2)(-4x+4) x^2(x-1)4(x-1) x^2+4 and x-1 x^2+4=0 x^2=-4 square root it and that's the answer which is +2i and -2i x-1 x=0,+ and -2i

OpenStudy (beginnersmind):

you copied the wrong equation hs. Anyway writing it as \ x^3-x^2 - 4(x-1) = 0 x^2(x-1) - 4(x-1) = 0 (x^2-4)(x-1) = 0 etc.

OpenStudy (anonymous):

thanks